# Examining Racial Stereotypes in YouTube Autocomplete Suggestions

### v1

@@ 1 Introduction
Most prior research on autocomplete suggestions centers on algorithmic outputs from general-purpose search engines like Google that can guide users to textual or image content (Al-Abbas et al., 2020; Baker and Potts, 2013; Leidinger and Rogers, 2023; Roy and Ayalon, 2020). This research focus seems reasonable because Google Search has been the largest global search engine since the early 2000’s. [^2] https://www.britannica.com/topic/Google-Inc However, the use of autocomplete to assist query inputs goes far beyond Google searching. For instance, YouTube, the platform we focus on in this study, has incorporated the autocomplete feature in its search function for many years. Indeed, if we count the number of searches made across all platforms, YouTube emerges as the second-largest search engine after Google (Smith, 2020), offering all its results in a video format. However, YouTube autocomplete has attracted little scholarly attention, perhaps because of YouTube’s public perception as a video-sharing platform rather than a search engine.
@@
Given YouTube’s wide reach, the cultural influence of its offerings, and the unique content curation mechanisms that shape its search outputs (including both search results and autocomplete suggestions), we argue that it is crucial to examine how YouTube’s search algorithms might affect end-users’ information gathering and belief formation processes. We begin to fill this vital research gap by exploring what information YouTube users in the United States may encounter through the autocomplete suggestions provided by YouTube’s search function. Since misinformation or limited information about racial groups, along with media stereotypes, can shape individuals’ perceptions of race (Covington, 2010), we pose the following research question: what information does YouTube provide to users exploring information about racial topics, and how may that information shape users’ understandings of race?
@@
Importantly, the autocomplete feature not only provides information but may also contribute to the propagation of racial stereotypes. We perform an algorithmic output audit of YouTube autocomplete suggestions regarding four racial groups—Whites, Blacks, Asians, and Hispanics—to surface such stereotypes. We contend that discriminatory race-related outputs generated by algorithms, including autocomplete suggestions, should be interrogated as the most visible signs of deeper issues related to racial biases inherent in sociotechnical systems driving such algorithms. Given that autocomplete outputs on race-related topics are shaped by the popularity of videos uploaded to YouTube, studying them also offers us insights into the broader sociocultural discourses on racial issues.
@@ 2.1 Racial Stereotypes
Asians or Asian Americans, often described as a ‘model minority,’ are stereotyped as being intelligent, hard-working, and academically and economically accomplished (Lee et al., 2008; Trytten et al., 2012). While this model minority stereotype could be perceived as positive, Lee et al. (2008) argued that it can conceal the diversity among Asians, rouse interracial tensions, and discourage Asians from disclosing their problems and seeking assistance. In addition to the model minority stereotype, Asians have a long history of being marginalized in the U.S. Chen (1996) showed that the physical appearance of Asian male immigrants in the nineteenth century and the work they engaged in were often caricatured to build feminized and infantilized images; the mass media at that time reinforced this image by portraying them as wearing feminine clothes or engaging in housework (Chen, 1996). Further, the U.S. mass media often portrayed Asian women as exotic, submissive, and quiet (Paner, 2018).
Another US minority, Hispanics, tend to be categorized as a single group despite their heterogeneous and complex ethnic roots (Nelson and Tienda, 2014). They were typically described in early Western films as highly emotional and violent individuals with untidy features, aka the ‘El Bandido’ stereotype (Berg, 2002). This stereotype has gradually evolved into the image of undocumented immigrants, contributing to racist discourses that perceive them as criminals who are potentially threatening to American lives and jobs (Pérez Huber and Solorzano, 2015).
@@ 2.2 Algorithmic Bias, Search Critiques, and Algorithm Audits
As digital platforms grow, systems designers rely on algorithms to determine what information to present to whom. These algorithms and the interfaces created for them can shape people’s behaviors and perceptions of the world (Kay et al., 2015). However, both algorithms and interfaces can exhibit biases in their representation of information. While there is no universally agreed-upon definition of algorithmic bias, researchers from various disciplines have addressed its different manifestations. In social science, scholars tend to focus on how algorithmic bias perpetuates social discrimination and affects social equity.
@@
Prior research has shown algorithmic biases in the operation of search engines regarding what they index, what they present to specific users, and—most relevant to our work—how their autocomplete suggestions reinforce social discrimination (Kay et al., 2015). For example, Baker and Potts (2013) showed that Google search produces a higher number of negatively stereotyped autocomplete suggestions for searchers with Muslim, Jewish, Gay, or Black identities. Lin et al. (2023) collected autocomplete predictions from three leading search engines—Google, Bing, and Baidu—and showed disparities in their toxicity scores with respect to gender, race, and sexual orientation.  Leidinger and Rogers (2023) showed that autocompletes from Yahoo, another general-purpose search engine, include a large number of negative stereotypes for Latinos, e.g., portraying them as stupid and loud. Our work contributes to this line of research on autocomplete algorithm audits and the measurement of algorithmic biases.
@@ 3.2 Data Collection
We collected data from July through August 2023 by gathering YouTube autocomplete suggestions for our input queries. Figure 2 shows an example of a YouTube search with a sample input query and its autocompletes; we collected such autocompletes for each query. We automated our data collection by developing Python scripts that simulated YouTube search querying actions. Specifically, we used Chrome as a web browser and accessed YouTube in incognito mode without logging in to any account to prevent the effects of personalization and browsing history. Our scripts opened and closed a new browser window for each query to ensure that the search history of other input queries did not influence the autocomplete suggestions for any single query.
@@
We used a virtual private network (VPN) to collect data from multiple locations in the U.S. to account for location-specific biases (Ribeiro et al., 2020). In particular, we collected YouTube autocompletes from five states: New Jersey, California, Georgia, Texas, and Washington. In addition, we conducted our data collection three times, each separated by a few days, to account for the temporal variations in algorithmic suggestions. Through these efforts, we attempted to build a geographically representative and temporally stable dataset of autocomplete suggestions across the U.S.
Analyzing the data samples from three different waves and five different states, we found minimal differences among them. We integrated these samples and removed duplicate results to create a combined dataset for further inquiry. This final dataset consists of 246 autocomplete results. We describe details of our data preparation process, descriptive statistics, and inter-sample differences in the Supplemental Material.
@@ 3.3 Data Analysis
We used inductive analysis (Elo and Kyngäs, 2008) to examine our data and extract concepts and categories that describe the underlying phenomena. Prior studies on investigating the biases and stereotypes in Google’s autocompletes have used a similar approach for data analysis (Al-Abbas et al., 2020; Roy and Ayalon, 2020); their findings showed the strength of this approach in analyzing and categorizing the biases underlying the autocompletes. We present additional details of our coding and analysis, including how we attended to video results for autocompletes, in the Supplemental Material.
Throughout the process of qualitative coding and categorization of autocomplete results, we (the authors) incorporated critical discourse analysis (CDA). CDA is a systematic approach to examining a text, event, or discursive practice to identify how it originates from and is affected by power relations (Fairclough, 1993). We employed this framework to analyze each autocomplete as concurrently a text, a discourse practice, and a sociocultural practice (Fairclough, 1995). Using this approach, we aimed to identify various facets of each discursive event, thereby revealing particular power dynamics reflected in that event. We engaged in regular discussions about how the autocomplete suggestions in our data reproduce and legitimize social inequalities, and our coding sought to identify various forms of domination and biases.
@@ 4 Results
Main category Subcategory Description
Appearance Personal Hygiene The activity of cleaning the body and maintaining a good appearance
Skin Tone The skin color originating from racially different genetic factors
Ability Talent One’s abilities for artistic or physical performances
Financial One’s abilities to purchase goods or services
Intellectual One’s abilities to learn and study
Culture Cultural Heritage Conventional cultural elements of one society with a shared identity
Ethnic Humor The humor based on racial, ethnic, or national stereotypes
Language How someone communicates verbally
Relationships How people interact and connect with each other
Social Equity Diversity/Inclusion Practices of embracing socially diverse groups of people
Racial Justice Pursuit of fighting against injustice toward a certain racial identity
Manner Aggression Antisocial and violent traits
Inappropriate behavior Being idle or behaving in a socially inept way
Table 2: Categories of racial stereotypes that emerged in YouTube autocomplete suggestions.
@@ 4.1 Appearance
Personal Hygiene. This category is notable because only Blacks were subject to autocompletes related to personal hygiene. These autocompletes referenced Black people washing their hair or beard (“How often should a Black man/woman wash his/her hair?,” “How often should a Black man wash his beard?”) and shaving (“Should a Black man shave every day?”). Figure 3333https://www.youtube.com/watch?v=CZCuI4x1sUc shows the preview of a video resulting from the suggested autocompletes, which contains advice for preventing white flakes in hair. Another autocomplete in this category asked, “Do Black people get lice?” It is possible that these autocompletes reflect highly sought-after searches by Black people for the information needed to take care of textured hair. However, these questions are raised only for the Black group in our data, raising the possibility that they reflect the racial stereotypes of Black people having poor hygiene. Historically, a deep-rooted racial stereotype maintained that dark or non-white skin was associated with uncleanness (Brown, 2009); indeed, this stereotype served as a rationale for racial segregation in the U.S. (Lintner, 2004). These YouTube autocompletes suggest that the negative stereotypes about Blacks’ hygiene unfortunately persist.
Figure 3: A YouTube video suggested by the autocomplete “How often should a Black man wash his hair?.”
Skin Tone. Similar to ‘Personal Hygiene,’ only Blacks were subjects of autocompletes in the ‘Skin’ category. Autocompletes related to skin tones included, “Can a Black person become fair?” and “Can a Black person bleach their skin?” These bleaching-related autocompletes reflect both an interest in lighter skin and the implicit negative view of darker skin (Hunter, 2011). We observed no autocompletes related to the action of changing natural skin tones for other races.
Other autocompletes show how Black skin is perceived as unique and distinct from other races. Examples include autocompletes asking about the need for Black people’s skin protection (“Should Black people wear sunscreen?,” “Do Black people get sunburn?”) and changes in their natural skin tones (“Can Black people blush?,” “Can Black people tan?”). These results suggest both an otherization and general lack of information about Black skin characteristics.
@@ 4.2 Ability
Talent. Our analysis revealed racial stereotypes associated with talent in music, dance, and sports that mainly addressed White and Black groups. First, for White groups, the autocompletes related to their artistic talent included both positive (“White girl can rap,” “White girl can sing”) and negative (“White people can’t dance”) elements. Regarding physical talent, several autocompletes reflect stereotyped perceptions that Whites are not as athletic as Blacks (“White men can’t jump,” “White boys can’t jump”). Indeed, the videos that such autocompletes guided users to related to a 1992 sports comedy movie called “White Men Can’t Jump,” which leverages the widely held stereotype that race is a positive indicator of physical ability (Felson, 1981). Autocomplete suggestions linked to this movie’s title decades after its release date suggest a cultural appetite to mock or stereotype White athletes as being inferior in their physical performance.
@@
Similarly, most autocompletes for Blacks referenced their artistic or physical ability in both positive (“Black people can dance to anything,” “Black man can sing,”) and negative (“Black man can’t dance,” “Black man can’t play basketball”) terms. Notably, we found some autocompletes related to Blacks’ exceptional athletic abilities (“Are Black people more athletic?”), which aligns with the previously reported stereotype (Biernat and Manis, 1994) that Blacks are athletically superior to Whites. Those who did not meet such expectations were negatively stereotyped. For example, the autocomplete “Black man can’t dance/play basketball” reflects a stereotype that Black people should be good at dancing or basketball naturally, and the resulting videos mock Blacks who are not good dancers or sportsmen. In sum, these findings show that YouTube autocompletes reflect racial stereotypes based on widely held beliefs about the relationships between race and physical abilities.
@@
Financial. Only Blacks were subjects of autocompletes that belonged to financial ability. All autocompletes here were associated with a lack of financial resources and acumen, which reinforces poor Blacks as the representative public image of poverty. Interestingly, this is consistent with how network TV news and weekly news magazines have historically grossly overrepresented Blacks in their portrayal of poor people as a whole  (Gilens, 1996). Autocompletes in this category promoted the stereotype that Blacks cannot purchase or own expensive products, such as fancy cars or clothes. For example, “Black man cannot own g wagon,” “Black woman can’t buy a dress,” and “A Black man can’t have a suitcase” belonged to this category.
@@
While these results superficially reinforce the long-standing stereotypes of Blacks as poor, some video results for these autocompletes counter this stereotype. For example, the top video result for “Black woman can’t buy a dress” portrays a White shop assistant and manager treating a Black woman attempting to buy a bridal gown at a shop differently than they treat a White woman buying a gown (Figure 4). [^4] https://www.youtube.com/watch?v=XMcRS2-la3E This video undermines the stereotypical attitudes toward Black people regarding their financial ability (it also critiques the assumptions of criminality and poor hygiene associated with Blacks). Similarly, resulting videos for other autocompletes dismiss the assumption that Black people are poor by explicitly showing that Black people can and do own expensive products. This shows that users, i.e., users who click and watch, have opportunities to watch video content that helps diminish negative stereotypes of Blacks regarding financial ability.
@@
Intellectual. Intelligence-related autocompletes appeared for Asians and Blacks in our data. In the U.S., Asians have historically been perceived as intelligent, hard-working, and pursuing high educational achievements (Trytten et al., 2012). Autocompletes such as “Are Asians smarter?” reflect this model minority stereotype. Unlike Asians, Blacks have often been negatively stereotyped regarding their intellectual and academic abilities (Lintner, 2004; Miller, 1995). In line with this, we found two autocompletes regarding Blacks’ intelligence: “Black people can’t name one African country” and “Black man can’t say beginning.” Videos for the first autocomplete show the lack of knowledge of Black people about Africa, which suggests their ignorance of Blacks’ historical or ethnic roots. The second leads to a video of a South African President, who is Black, repeatedly mispronouncing his words when attempting to say “in the beginning.” These cases show how videos uploaded to YouTube position individual mishaps to denigrate racial groups under the guise of humor. YouTube’s automated use of such content to build autocomplete suggestions further removes necessary context and can thus perpetuate racial stereotypes.

### v2

@@ 1 Introduction
Most prior research on autocomplete suggestions centers on algorithmic outputs from general-purpose search engines, especially Google Search (Al-Abbas et al., 2020; Baker and Potts, 2013; Leidinger and Rogers, 2023; Roy and Ayalon, 2020). This research focus seems reasonable because Google Search has been the largest global search engine since the early 2000’s. [^2] https://www.britannica.com/topic/Google-Inc However, the use of autocomplete to assist query inputs goes far beyond Google searching. For instance, YouTube, the platform we focus on in this study, has incorporated the autocomplete feature in its search function for many years. Indeed, if we count the number of searches made across all platforms, YouTube emerges as the second-largest search engine after Google (Smith, 2020), offering all its results in a video format. However, YouTube autocomplete has attracted little scholarly attention, perhaps because of YouTube’s public perception as a video-sharing platform rather than a search engine.
Unlike search engines, YouTube allows users to not just view, but also create content on its site (Wotanis and McMillan, 2014). Prior research on YouTube’s participatory culture has characterized it as a critical form of expression, “self-promotion, escapism, and utter play.” (Boxman-Shabtai, 2019) However, it is possible that YouTube’s complex cultural fabric comprising both professional (i.e., traditional media) and amateur (i.e., independent individuals) content production (Boxman-Shabtai, 2019) induces the types of biases that do not appear on search engines.
YouTube algorithms construct and shape content creators’ visibility on the site’s search and front pages as well as in features like autoplaying of ‘related videos’ (Bishop, 2019). However, YouTube releases limited information about these algorithms to maintain a competitive advantage. This black-boxing of algorithmic governance makes accountability difficult (Reynolds and Hallinan, 2024). Thus, there are growing concerns about how cultural producers, who must navigate the increasingly algorithmic nature of content distribution on YouTube, risk losing their visibility and income (Bishop, 2019). Given YouTube’s unique participatory culture and the uneven power relationship between content creators or searchers and the platform, Reynolds and Hallinan (2024) have called for examining YouTube algorithms’ accountability in its specific platform environment. Responding to this call, we focus on YouTube’s search algorithms.
@@
Specifically, we explore what information YouTube users in the United States (U.S.) encounter through the search autocomplete suggestions. Since misinformation or limited information about racial groups, along with media stereotypes, can shape individuals’ perceptions of race (Covington, 2010), we pose the following research question: what information does YouTube provide to users exploring information about racial topics, and how may that information shape users’ understandings of race? We focus on race here to extend prior conversations on the media’s construction of race (Chen, 1996; Squires, 2014).
@@
Importantly, the autocomplete feature not only provides information but may also contribute to the propagation of racial stereotypes. We perform an algorithmic output audit of YouTube autocomplete suggestions regarding four racial groups—Whites, Blacks, Asians, and Hispanics—to surface such stereotypes. We contend that discriminatory race-related outputs generated by algorithms, including autocomplete suggestions, should be interrogated as the most visible signs of deeper issues related to racial biases inherent in sociotechnical systems driving such algorithms.
@@ 2.1 Racial Stereotypes
Asians or Asian Americans, often described as a ‘model minority,’ are stereotyped as being intelligent, hard-working, and academically and economically accomplished (Lee et al., 2008; Trytten et al., 2012). While this model minority stereotype could be perceived as positive, Lee et al. (2008) argued that it can conceal the diversity among Asians, rouse interracial tensions, and discourage Asians from disclosing their problems and seeking assistance. In addition to the model minority stereotype, Asians have a long history of being marginalized in the U.S. Chen (1996) showed that the physical appearance of Asian male immigrants in the nineteenth century and the work they engaged in were often caricatured to build feminized and infantilized images; the mass media at that time reinforced this image by portraying them as wearing feminine clothes or engaging in housework (Chen, 1996). Further, the U.S. mass media has often portrayed Asian women as exotic, submissive, and quiet (Paner, 2018).
Another US minority, Hispanics, tend to be categorized as a single group despite their heterogeneous and complex ethnic roots (Nelson and Tienda, 2014). They were typically described in early Western films as highly emotional and violent individuals with untidy features, aka the ‘El Bandido’ stereotype (Berg, 2002). This stereotype has gradually evolved into the image of undocumented immigrants, contributing to racist discourses that portray them as criminals who are potentially threatening to American lives and jobs (Pérez Huber and Solorzano, 2015).
@@ 2.2 Algorithmic Bias, Search Critiques, and Algorithm Audits
As digital platforms grow, systems designers rely on algorithms to determine what information to present to whom. These algorithms and the interfaces created for them can signal the quality and relevance of different search results and shape people’s perceptions of topics of societal importance (Kay et al., 2015). However, both algorithms and interfaces can exhibit biases in their representation of information. While there is no universally agreed-upon definition of algorithmic bias, researchers from various disciplines have addressed its different manifestations. In social science, scholars tend to focus on how algorithmic bias perpetuates social discrimination and affects social equity.
@@
Prior research has shown algorithmic biases in the operation of search engines regarding what they index, what they present to specific users, and—most relevant to our work—how their autocomplete suggestions reinforce social discrimination (Kay et al., 2015). For example, Baker and Potts (2013) showed that Google search produces a higher number of negatively stereotyped autocomplete suggestions for searchers with Muslim, Jewish, Gay, or Black identities. Lin et al. (2023) collected autocomplete predictions from three leading search engines—Google, Bing, and Baidu—and showed disparities in their toxicity scores with respect to gender, race, and sexual orientation.  Leidinger and Rogers (2023) showed that autocompletes from Yahoo, another general-purpose search engine, include a large number of negative stereotypes for Latinos, e.g., portraying them as stupid and loud.
Our work contributes to this line of research by deploying a novel methodological approach and conceptual focus to understanding the prevalence of racial biases in YouTube search autocompletes. While prior studies have primarily focused on only the text of autocomplete suggestions (Al-Abbas et al., 2020; Baker and Potts, 2013; Leidinger and Rogers, 2023; Roy and Ayalon, 2020), we analyze autocomplete texts in conjunction with the video results for each suggestion. This helps us present a more nuanced understanding of how YouTube users’ interactions with its search feature may influence their perspectives. Our work is the first to deploy critical discourse analysis (CDA) in addition to inductive analysis to uncover not only how language and discourse practices of autocompletes reflect problematic socio-cultural norms but also how they reinforce hegemonic whiteness and racist ideologies.
It is possible that biased queries entered by users, in turn, could be influencing the autocomplete algorithm to produce stereotypical outputs. However, given the black-box nature of autocomplete algorithm, the extent to which this happens remains unclear. Some recent evidence suggests that the impact of search volume on autocomplete suggestions may be lower than is generally assumed and, in some instances, prior user queries may not feature at all in certain suggestions (Graham, 2022). Thus, we largely focus in this article on the audit of autocomplete outputs rather than on inferring how biased user queries could influence the production of stereotypes in autocompletes.
@@ 3.2 Data Collection
We collected data from July through August 2023 by gathering YouTube autocomplete suggestions for our input queries. Figure 2 shows an example of a YouTube search with a sample input query and its autocompletes; we collected such autocompletes for each query. We automated our data collection by developing Python scripts that simulated YouTube search querying actions. This automated process comprised three stages: 1) accessing the YouTube.com website, 2) entering an input query in the YouTube search box, and 3) collecting autocomplete results, if available, and storing them. We used Chrome as a web browser and accessed YouTube in incognito mode without logging in to any account to prevent the effects of personalization and browsing history. Our scripts opened and closed a new browser window for each query to ensure that the search history of other input queries did not influence the autocomplete suggestions for any single query.
@@
We used a virtual private network to collect data from multiple locations in the U.S. to account for location-specific biases (Ribeiro et al., 2020). In particular, we collected YouTube autocompletes from five states: New Jersey, California, Georgia, Texas, and Washington. We conducted our data collection three times, each separated by a few days, to account for the temporal variations in algorithmic suggestions. Through these efforts, we attempted to build a geographically representative and temporally stable dataset of autocomplete suggestions across the U.S.
Analyzing the data samples from three different waves and five different states, we found minimal differences among them. We integrated these samples and removed duplicate results to create a combined dataset for further inquiry. While this step precluded prioritizing more geographically and temporally stable autocompletes, it allowed us to work with a more comprehensive set of suggestions. Examining the resulting dataset, we also noticed that some autocomplete suggestions were not related to the racial groups. Thus, we further removed the irrelevant autocomplete suggestions based on whether the autocomplete refers to non-human rather than non-human subjects. After removing duplicates, our dataset consisted of a total of 241 unique autocompletes.
We describe more details of our data preparation process and descriptive statistics in the Supplemental Material.
@@ 3.3 Data Analysis
We used inductive analysis (Elo and Kyngäs, 2008) to examine our data and extract concepts and categories that describe the underlying phenomena. Two authors conducted a bottom-up open coding process on the platform ‘Dedoose’ which allowed the authors to collaborate in real-time.
We began by developing a situated understanding of each autocomplete suggestion by examining about ten videos displayed as top search results when selecting that suggestion. This examination helped us recognize the contextual and cultural meaning of autocomplete suggestions and we leveraged this nuanced understanding to develop our open codes. This careful, context-driven coding process was important because some of the autocomplete results were closely related to memes or video content in their corresponding search results. For instance, the autocomplete “when an asian is kidnapped” was related to comedy content created by an Asian YouTuber. Thus, it was important to search for and understand the situated meanings of autocompletes rather than interpret them in isolation.
Then, we developed the common themes arising from our open codes and their corresponding autocomplete data by grouping similar codes into conceptual categories. At this point in the process, when any pre-existing categories seemed not to directly align with an emergent theme, we carefully considered whether to introduce a new category, revise an existing category, and/or create a subcategory. This process was iterative and we compared our emerging categories with one another and with the underlying data at each stage of our analysis. Some autocompletes represented miscellaneous sociocultural contexts that did not warrant inclusion in or creation of a separate category; we excluded such autocompletes from further analysis. In total, we grouped 217 autocompletes into a two-tier category structure.
Throughout the process of qualitative coding and categorization of autocomplete results, we incorporated critical discourse analysis (CDA). CDA is a systematic approach to examining a text, event, or discursive practice to identify how it originates from and is affected by power relations (Fairclough, 1993). We employed this framework to analyze each autocomplete as concurrently a text (language produced in a discursive event), a discourse practice (the production, distribution, and consumption of a text), and a sociocultural practice (hegemonic struggles that ideologically shape a discursive practice) (Fairclough, 1995). This meant that during data analysis, our focus was less on examining the lexical meanings of each autocomplete suggestion than on the hegemonic function the suggestion performed or indicated within the sociocultural practices represented by our five main categories. Using this approach, we aimed to identify various facets of each discursive event, thereby revealing particular power dynamics reflected in that event. We engaged in regular discussions about how the autocomplete suggestions in our data reproduce and legitimize social inequalities, and our coding sought to identify various forms of domination and biases.
@@ 4 Results
Main category Subcategory Description % (n)
Appearance Personal Hygiene The activity of cleaning the body and maintaining a good appearance 3 (6)
Skin Tone The skin color originating from racially different genetic factors 5 (10)
Ability Talent One’s abilities for artistic or physical performances 22 (47)
Financial One’s abilities to purchase goods or services 2 (4)
Intellectual One’s abilities to learn and study 2 (4)
Culture Cultural Heritage Conventional cultural elements of one society with a shared identity 21 (45)
Ethnic Humor The humor based on racial, ethnic, or national stereotypes 13 (28)
Language How someone communicates verbally 6 (12)
Relationships How people interact and connect with each other 9 (19)
Social Equity Diversity/Inclusion Practices of embracing socially diverse groups of people 2 (5)
Racial Justice Pursuit of fighting against injustice toward a certain racial identity 10 (22)
Manner Aggression Antisocial and violent traits 5 (11)
Inappropriate Behavior Being idle or behaving in a socially inept way 2 (4)
Table 2: Categories of racial stereotypes that emerged in YouTube autocomplete suggestions. % and n refer to the percentage and number of autocompletes belonging to each subcategory in the analyzed data, respectively.
@@ 4.1 Appearance
Personal Hygiene. This category is notable because only Blacks were subject to autocompletes related to personal hygiene. These autocompletes referenced Black people washing their hair or beard (“how often should a black man wash his hair,” “how often should a black woman wash her hair,” “how often should a black man wash his beard”) and shaving (“should a black man shave everyday”). Figure 3 shows the preview of a video resulting from the suggested autocompletes, which contains advice for preventing white flakes in hair. [^3] https://www.youtube.com/watch?v=CZCuI4x1sUc Another autocomplete asked, “do black people get lice.” It is possible that these autocompletes reflect highly sought-after searches by Black people for the information needed to take care of textured hair. However, these questions are raised only for the Black group, raising the possibility that they reflect the racial stereotypes of Black people having poor hygiene. Historically, a deep-rooted racial stereotype maintained that dark or non-white skin was associated with uncleanness (Brown, 2009); indeed, this stereotype served as a rationale for racial segregation in the U.S. (Lintner, 2004). These YouTube autocompletes suggest that the negative stereotypes about Blacks’ hygiene unfortunately persist.
Figure 3: A YouTube video suggested by the autocomplete “how often should a black man wash his hair.”
Skin Tone. Similar to ‘Personal Hygiene,’ Blacks were the main subjects of autocompletes in the ‘Skin Tone’ category. Autocompletes related to skin tones included “can a black person become fair.” These bleaching-related autocompletes reflect both an interest in lighter skin and the implicit negative view of darker skin (Hunter, 2011). We observed no autocompletes related to the action of changing natural skin tones for other races. Other autocompletes show how Black skin is perceived as unique and distinct from other races. Examples include autocompletes asking about the need for Black people’s skin protection (“should black people wear sunscreen,” “do black people get sunburn”) and changes in their natural skin tones (“can black people blush,” “can black people tan”). These results suggest both an otherization and a general lack of awareness about Black skin characteristics.
@@ 4.2 Ability
Talent. Our analysis revealed racial stereotypes associated with talent in music, dance, and sports that mainly addressed White and Black groups. For White groups, the autocompletes related to their artistic talent included both positive (“white girl can rap,” “white girl can sing”) and negative (“white people can’t dance”) elements. Regarding physical talent, several autocompletes reflect stereotyped perceptions that Whites are not as athletic as Blacks (“white men can’t jump,” “white boys can’t dance”). Indeed, the videos that such autocompletes guided users to related to a 1992 sports comedy movie called “White Men Can’t Jump,” which leverages the widely held stereotype that Black race is a positive indicator of physical ability (Felson, 1981). Autocomplete suggestions linked to this movie’s title decades after its release date suggest a cultural appetite to mock or stereotype White athletes as being inferior in their physical performance.
@@
Similarly, most autocompletes for Blacks referenced their artistic or physical ability in both positive (“black people can dance to anything,” “black man can sing”) and negative (“black man can’t dance,” “black man can’t play basketball”) terms. Notably, we found some autocompletes related to Blacks’ exceptional athletic abilities (“are black people more athletic”), which aligns with the previously reported stereotype (Biernat and Manis, 1994) that Blacks are athletically superior to Whites. Those who did not meet such expectations were negatively stereotyped. For example, the autocompletes “black man can’t dance,” or “black man can’t play basketball” reflect a stereotype that Black people should be good at dancing or basketball naturally, and the resulting videos mock Blacks who are not good dancers or sportsmen. In sum, these findings show that YouTube autocompletes reflect racial stereotypes based on widely held beliefs about the relationships between race and physical abilities.
@@
Financial. Only Blacks were subjects of autocompletes that belonged to financial ability. All autocompletes here were associated with a lack of financial resources and acumen, which reinforces poor Blacks as the representative public image of poverty. Interestingly, this is consistent with how network TV news and weekly news magazines have historically grossly overrepresented Blacks in their portrayal of poor people as a whole (Gilens, 1996). Autocompletes in this category promoted the stereotype that Blacks cannot purchase or own expensive products, such as fancy cars or clothes (“black man cannot own g wagon,” “black woman can’t buy a dress,” “a black man can’t have a suitcase”).
@@
While these results superficially reinforce the long-standing stereotypes of Blacks as poor, some video results for these autocompletes counter this stereotype. For example, the top video result for “black woman can’t buy a dress” portrays a White shop assistant and manager treating a Black woman attempting to buy a bridal gown at a shop differently than they treat a White woman buying a gown (Figure 4). [^4] https://www.youtube.com/watch?v=XMcRS2-la3E This video undermines the stereotypical attitudes toward Black people regarding their financial ability (it also critiques the assumptions of criminality and poor hygiene associated with Blacks). Similarly, resulting videos for other autocompletes about financial ability dismiss the assumption that Black people are poor by explicitly showing that Black people can and do own expensive products. This shows that in some cases, users who review and watch the search results may get exposed to video content that helps diminish negative stereotypes of Blacks regarding financial ability.
@@
Intellectual. Intelligence-related autocompletes appeared for Asians and Blacks. In the U.S., Asians have historically been perceived as intelligent, hard-working, and pursuing high educational achievements (Trytten et al., 2012). Autocompletes such as “are asians smarter” reflect this model minority stereotype. Unlike Asians, Blacks have often been negatively stereotyped regarding their intellectual and academic abilities (Lintner, 2004; Miller, 1995). In line with this, we found autocompletes regarding Blacks’ intelligence: “black people can’t name one african country,” and “black man can’t say beginning.” Videos for the first autocomplete show the lack of knowledge of some Black people about Africa, which suggests their ignorance of Blacks’ historical or ethnic roots. The second leads to a video of a South African President, who is Black, repeatedly mispronouncing his words when attempting to say “in the beginning.” These cases show how videos uploaded to YouTube position individual mishaps to denigrate racial groups under the guise of humor. YouTube’s automated use of such content to build autocomplete suggestions further removes necessary context and can thus perpetuate racial stereotypes.